Solve the problems below and show the complete solution.

A vertical spring has a force constant k of 130N/m and a mass of 0.645kg is placed on top of it.
a) Calculate the mass of the object in N. F = w g
b) Calculate the x if the spring be compressed. F = k x
x = F / k
1
Expert's answer
2020-12-10T08:48:40-0500

a) We can find the mass of the object from the formula:


"F=mg=0.645\ kg\cdot 9.8\ \dfrac{m}{s^2}=6.32\ N."

b) We can calculate "x" from the Hooke's Law:


"F=kx,""x=\dfrac{F}{k}=\dfrac{6.32\ N}{130\ \dfrac{N}{m}}=0.048\ m."

Answer:

a) "F=6.32\ N."

b) "x=0.048\ m."